# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

### v2

@@ 7 Conclusion
The present paper proposes a novel personalized federated learning framework that integrates heat-kernel enhanced multi-view clustering with tensor-based global coordination. The primary contributions of this study include the introduction of a novel method, Heat-kernel Enhanced Federated Clustering, which has not been previously documented. The adoption of heat-kernel coefficients in federated multi-view clustering facilitates the more effective capture of complex local patterns while ensuring the maintenance of global coordination. In the Personalized Aggregation Mechanism, the framework provides an adaptive personalization that balances local specialization with global knowledge sharing through learnable personalization parameters. In the context of Privacy-Preserving Protocol, the aggregation protocol has been engineered to ensure differential privacy while preserving the efficacy of clustering through the implementation of meticulously designed statistical sharing methodologies. A comprehensive theoretical analysis is provided, offering convergence guarantees, privacy bounds, and complexity analysis. These elements serve to establish the theoretical foundation of the approach.
@@ 7.1 Future Directions
This work suggests several promising research directions. Initially, the Dynamic View Discovery method was employed. The framework is being extended to accommodate the dynamic alteration of view structures during federated rounds. The present study explores the potential of hierarchical federated clustering for large-scale deployment, investigating multi-level federated architectures in the process. In the context of adversarial robustness, the focus is on the analysis and enhancement of systems’ resilience to malicious clients and data poisoning attacks. The concept of continuous learning involves the adaptation of a framework for federated continual learning scenarios, with the understanding that data distributions undergo evolution over time. In the domain of cross-modal applications, the focus is on the exploration of applications to federated learning with fundamentally different data modalities, including but not limited to text, images, and sensors. The proposed framework presents novel opportunities for privacy-preserving, personalized machine learning on distributed multi-view data, with extensive applications in healthcare, the Internet of Things (IoT), and collaborative intelligence systems. The proposed framework opens new possibilities for privacy-preserving, personalized machine learning on distributed multi-view data, with broad applications in healthcare, IoT, and collaborative intelligence systems.
@@ 7.2 Potential Limitations
The proposed frameworks, while innovative, present several potential limitations. The computational complexity associated with tensor decompositions may pose challenges for clients with limited resources. The selection of personalization parameters requires careful tuning to balance local adaptation and global coordination effectively. The assumption of data homogeneity across clients may not hold in all scenarios, potentially impacting clustering performance. Additionally, the privacy guarantees depend on the proper calibration of noise, which may affect the utility of the shared statistics. Future work will focus on addressing these limitations through algorithmic optimizations, adaptive parameter selection methods, and robust privacy-preserving techniques with minimal utility loss and enhanced adaptability to heterogeneous data distributions. Validations both on synthetic and real-world datasets will be conducted to assess the practical performance and limitations of the proposed frameworks.
@@ 7.3 Applications and Implications
The proposed personalized federated heat-kernel enhanced multi-view clustering framework has significant applications across various domains. In healthcare, it can facilitate collaborative analysis of patient data from multiple institutions while preserving privacy. In IoT, the framework can enable efficient clustering of sensor data collected from distributed devices, enhancing real-time decision-making. In collaborative intelligence systems, it can support the integration of diverse data sources for improved insights and predictions. The implications of this work extend to advancing privacy-preserving machine learning techniques, promoting personalized models that respect individual client characteristics, and fostering collaboration in distributed environments without compromising data security. With the aggregation of several factors such as privacy, personalization, and multi-view data integration using tensorization, this framework sets a new standard for federated learning applications in complex, real-world scenarios.
Acknowledgment: This paper was authored during the period of the author’s postdoctoral studies at ISTI-CNR, Italy.

### v3

@@ 5 Conclusion
This paper introduces innovative personalized federated learning frameworks that integrate heat-kernel enhanced multi-view clustering with tensor-based global coordination, PARAFAC2, and Tucker decompositions. The introduction is accompanied by guarantees, privacy bounds, and complexity analysis. These elements serve to establish the theoretical foundation of the proposed approaches. The incorporation of heat-kernel coefficients in federated multi-view clustering facilitates the more effective capture of complex local patterns while ensuring the maintenance of global coordination. In the Personalized Aggregation Mechanism, the framework provides an adaptive personalization that balances local specialization with global knowledge sharing through learnable personalization parameters. In the context of Privacy-Preserving Protocol, the aggregation protocol has been engineered to ensure differential privacy while preserving the efficacy of clustering through the implementation of meticulously designed statistical sharing methodologies. The proposed frameworks effectively address the challenges associated with high-dimensional multi-view data in federated environments, ranging from matrix to tensor representations. The theoretical analysis encompasses convergence properties, privacy guarantees, and computational complexity, thereby providing a comprehensive understanding of the proposed methods’ performance and limitations.
@@ 5.1 Future Directions
This work suggests several promising research directions. The initial implementation of the dynamic view discovery method involved the utilization of weighting. The framework is being extended to accommodate the dynamic alteration of view structures during federated rounds. The present study explores the potential of hierarchical federated clustering for large-scale deployment, investigating multi-level federated architectures in the process. In the context of adversarial robustness, the focus is on the analysis and enhancement of systems’ resilience to malicious clients and data poisoning attacks. The concept of continuous learning involves the adaptation of a framework for federated continual learning scenarios, with the understanding that data distributions undergo evolution over time. In the domain of cross-modal applications, the focus is on the exploration of applications to federated learning with fundamentally different data modalities, including but not limited to text, images, and sensors. The proposed frameworks also present novel opportunities for privacy-preserving, customized machine learning on large language models (LLMs) and emerging multimodal models in federated settings. Subsequent research endeavors will investigate the integration of heat-kernel enhanced multi-view clustering with LLMs to facilitate collaborative learning from distributed text and multimodal data while ensuring privacy preservation. This encompasses the development of techniques for efficient model updates, the management of heterogeneous data distributions, and the assurance of robust privacy guarantees.
@@ 5.2 Potential Limitations
The proposed frameworks, while innovative, present several potential limitations. The computational complexity associated with tensor decompositions may pose challenges for clients with limited resources. The selection of personalization parameters necessitates meticulous calibration to harmonize local adaptation and global coordination with optimal efficacy. The assumption of data homogeneity across clients may not hold in all scenarios, potentially impacting clustering performance. Additionally, the efficacy of privacy guarantees is contingent upon the accurate calibration of noise, a process that has the potential to influence the utility of shared statistics. Subsequent endeavors will concentrate on ameliorating these limitations through the implementation of algorithmic optimizations, adaptive parameter selection methodologies, and robust privacy-preserving techniques that exhibit minimal utility degradation and augmented adaptability to heterogeneous data distributions. In order to evaluate the practical performance and limitations of the proposed frameworks, it is necessary to conduct validations on both synthetic and real-world datasets.
@@ 5.3 Applications and Implications
The proposed personalized efficient federated kernel multi-view clustering using tensor framework has significant applications across various domains. In the domain of healthcare, this technology has the potential to enable collaborative analysis of patient data from multiple institutions while ensuring the confidentiality of personal information. In the context of the Internet of Things (IoT), the framework has the potential to facilitate the effective aggregation of sensor data collected from distributed devices, thereby enhancing the efficacy of real-time decision-making processes. In the context of future smart cities, the proposed framework has the potential to facilitate the integration of multi-view data from diverse sources, including traffic sensors and social media platforms. This integration has the potential to enhance urban planning and service delivery, thereby promoting the development of more efficient and effective urban infrastructures. In the context of prospective industrial applications, the system has the capacity to enhance operational efficiency by means of data aggregation from disparate phases of the production process. In the context of future internet applications, this technology has the potential to enhance user experience by clustering multi-view data from various online platforms. In the context of collaborative intelligence systems, the framework has the potential to facilitate joint learning from heterogeneous data sources while ensuring the preservation of data ownership rights. The implications of this work extend to the advancement of privacy-preserving machine learning techniques, the promotion of personalized models in federated settings, and the fostering of collaboration across distributed data sources without compromising individual privacy. In light of the mounting prevalence of distributed data, the proposed framework provides a robust solution for multi-view clustering that achieves a balance between local adaptation and global knowledge sharing. This development signifies a significant advancement, paving the way for more effective and privacy-conscious machine learning applications.                                                                                                                                  Acknowledgment: The authorship of this paper is attributed to the period during which the author pursued postdoctoral studies at ISTI-CNR in Italy.